Circular No. 8108

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003da AND 2003db Additional apparent supernovae found by T. Boles (cf. IAUC 8103):

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003da Apr. 6.90 9^{h}23^{m}47.^{s}56 +42^{o}10′48′′.6 15.9 1′′.8 W, 12′′.0 S
2003db Apr. 7.00 9 36 58.83 +31 20 01.7 17.5 8′′.8 E, 9′′.4 N

Additional unfiltered CCD magnitudes from Boles’ images: SN 2003da in UGC 4992, Feb. 2, [19.0; Mar. 2, [19.0?; Apr. 7.83, 15.9. SN 2003db in MCG +05-23-21, Feb. 22, [19.0; Apr. 7.84, 17.5. Neither object is present on Palomar Sky Survey red or blue plates.

GRB 030329 P. Garnavich, University of Notre Dame; T. Matheson, Harvard-

Smithsonian Center for Astrophysics (CfA); D. Eisenstein, University of Arizona (UA); B. Pindor, Princeton University; N. Hathi, R. Jansen, R. Windhorst, and L. Echevarria, Arizona State University; J. Lee (UA); K. Krisciunas, Observatories of the Carnegie Institution of Washington (OCIW) and Cerro Tololo Interamerican Observatory; P. Martini, OCIW; and W. Brown, N. Caldwell, P. Berlind, M. Calkins, and K. Z. Stanek (CfA) report that spectra of the afterglow of GRB 030329 (IAUC 8101) have been obtained with the 6.5-m MMT, 6.5-m Magellan, and 1.5-m Tillinghast telescopes regularly from Mar. 30.12 to Apr. 7.25 UT. The early spectra consist of a power-law continuum with narrow emission lines originating from H II regions in the host galaxy at a redshift of z = 0.168. However, the spectra taken on Apr. 6 (and confirmed on Apr. 7) now show broad peaks in flux that are characteristic of a supernova. The broad bumps are seen at ≈ 570 and 480 nm, and these are similar to those in the spectrum of the peculiar type-Ic supernova 1998bw a week before maximum light (Patat et al. 2001, Ap.J. 555, 900). SN 1998bw was associated with GRB 980425 (IAUC 6884, 6895, 6896, 6899, 6901, 6903; Galama et al. 1998, Nature 395,

670) but the energy of that burst was 100 times less than that of classical γray-bursts (GRBs). The fading power-law spectrum suggests the presence of non-thermal emission, while broad features in the recent spectrum are a signature of atomic absorption from high velocity gas that is indicative of a supernova. The spectral similarity to SN 1998bw and other ‘hypernovae’ such as 1997ef (IAUC 6783, 6798, 6820; Iwamoto et al. 2000, Ap.J. 534,

660) provides strong evidence that at least some classical GRBs originate from core-collapse supernovae.

2003 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green